# Things we can learn by considering random locally symmetric manifolds

### v1

@@ 10 Stationary random subgroups
The notion of invariant random subgroups has proven very effective to the study of certain type of problems. However, the restriction to ‘invariant’ measures limits the scope of problems that one can investigate, in particular as the groups we are studying are non-amenable. For example, the rigidity theorem of Stuck and Zimmer says that for higher rank semisimple Lie groups (with property $(T)$) all IRS comes from lattices. In particular one cannot hope to study thin subgroups or general discrete subgroups which are non-lattices using the theory of IRS. Also for simple groups of rank one, it is known that there are discrete subgroups which cannot be in the support of an IRS. On the other hand every discrete subgroup is associated with some stationary random subgroups (hereafter SRS) as we shall explain below.
@@
Let $G$ be a connected semisimple Lie groups. Let $\mu$ be a Borel regular probability measure on $G$ whose support generates a Zariski dense subgroup. Let $X$ be a compact $G$-space. A probability measure $\nu\in\text{Prob}(X)$ is called $\mu$-stationary if $\mu*\nu=\nu$ where
@@
We let $(B,\nu)$ denote the Poisson boundary of $G$ associated to $\mu$. Then $\nu$ is $\mu$-stationary and has the following universal property. Recall that on a compact convex space every probability measure has a unique barycenter. In particular this applies to probability measure on the compact convex space $\text{Prob}(X)$.
@@
Let $X$ be a $G$-space. A probability measure $\nu_{0}$ on $X$ is $\mu$-stationary iff it is the $\nu$-barycenter of the measure $f_{*}\nu$ for some measurable map $f:B\to\text{Prob}(X)$. Moreover, the map $f$ is unique.
@@
Let $K$ be a fixed maximal compact subgroup of $G$. Below we will only consider measures $\mu$ which are bi-$K$-invariant. For such measures the Poisson boundary is $(G/P,\nu)$ where $P$ is a minimal parabolic subgroup and $\nu$ is the unique $K$-invariant measure on $G/P$ (the uniqueness is due to the fact that $K$ acts transitively on $G/P$ which is a consequence of the Iwasawa decomposition) . By the observation above we may vary the measure $\mu$ within this class (acording to our needs) without effecting the property of being stationary.
@@
Motivated by [Ge18b] and the original paper of Kazhdan and Margulis [KM68] we have establish in [GLM22] the following result which is a strong effective version of the Kazhdan–Margulis theorem. It is obviously stronger than the result from [Ge18b] as it is effective and applies to all SRS rather than IRS. Moreover, since it applies to all SRS it implies the Kazhdan–Margulis theorem for all discrete subgroups.
@@
$\nu(\{\Lambda\>:\>\Lambda\cap\mathrm{B}(r)\neq\{\mathrm{id}_{G}\}\})\leq\beta r^{\delta}\quad\forall r<\rho.$ (4)
As a straight forward consequence we obtain:
@@ 11 Confined subgroups have finite co-volume
In this section we will explain the proof of the following result from [FG23] which confirmed a conjecture of Margulis.
@@
Geometrically, the theorem says that every the random walk on a locally $X$-manifold of infinite volume, spans most of the time in the $r$-thick part for every $r>0$.
@@
The aim is to prove that $\nu_{n}\to\delta_{\langle 1\rangle}$. Since $\text{Prob}(\text{Sub}(G))$ (with the week-* topology) is compact, it is enough to prove that $\delta_{\langle 1\rangle}$ is the only limit of $\nu_{n}$. Thus let $\nu_{\infty}$ be a week-* limit of $\nu_{n}$, and we shall argue to show that $\nu_{\infty}=\delta_{\langle 1\rangle}$. In view of Theorem 10.5 $\nu_{\infty}$ is a discrete SRS, i.e. a $\nu_{\infty}$-random subgroup is almost surely discrete.
@@
Fix a minimal parabolic subgroup $P\leq G$ and recall that the Poisson boundary of $(G,\mu)$ is $(G/P,\nu_{P})$ where $\nu_{P}$ is the unique $K$-invariant probability measure on $G/P$. Let $\nu_{0}$ be an SRS. In view of Furstenberg’s theorem (see 10.1) $\nu_{0}$ is the baricenter of a measurable map $\phi:G/P\to\text{Prob}(\text{Sub}(G))$, that is:
@@
Let $G$ be a higher rank simple Lie group. Let $\mu$ be a smooth probability measure on $G$ and let $(X,\nu)$ be a probability $\mu$-stationary action of $G$. Then either $\nu$ is $G$-invariant, or there exists a proper parabolic subgroup $Q\subset G$ and a measure preserving $G$-equivariant map666 By measure preserving we only mean that $\nu(\pi^{-1}(A))=\mu_{Q}(A)$ for every Borel set $A\subset G/Q$. In particular we do not require that it is a relatively measure preserving extension. $\pi\colon(X,\nu)\to(G/Q,\mu_{Q}),$ where $\mu_{Q}$ is the unique $\mu$-stationary measure on $G/Q$.
@@
Arguing by way of contradiction, suppose that $\nu_{0}$ is a non-invariant discrete SRS. It follows from Theorem 11.4 that there is a parabolic subgroup $Q$ containing $P$ such that for $\nu_{P}$ almost all $gP\in G/P$ the measure $\phi(gP)$ is a $gPg^{-1}$-invariant discrete random subgroup of $gQg^{-1}$. The contradiction, thus follow from the following:
@@
Let $P\leq Q$ be parabolic subgroups of $G$. The only $P$-invariant discrete random subgroup of $Q$ is $\delta_{\langle 1\rangle}$.
@@
Let $Q$ be a parabolic subgroup of $G$. Let $Q=LN$ be a Levi decomposition of $Q$ and let $A_{L}$ be the centre of $L$. Then, any discrete $A_{L}$-invariant random subgroup of $Q$ is contained in $L$ almost surely.
@@
Let $\lambda$ be a discrete $A_{L}$-invariant random subgroup of $Q$. Suppose by way of contradiction that $\lambda(\{\Gamma\subset L\})<1$. Then there is a bounded identity neighbourhood $V_{2}$ in $Q$ such that $\lambda(\Omega)>0$ where $\Omega:=\{\Gamma:\Gamma\cap V_{2}\setminus L\neq\emptyset\}$. Furthermore, we may take $V_{2}$ to be of the form $V_{2}=W\cdot U_{2}$ where $W\subset L$ and $U_{2}\subset N$ and suppose that they are preserved by $a^{-1}$ (e.g. we can suppose that $\log U_{2}$ is a norm ball in the Lie algebra of $N$). Since $\lambda$ is discrete we can chose a small identity neighborhood $U_{1}\subset N$ so that setting $V_{1}=W\cdot U_{1}$ we have
@@
We remark that the analog of stiffness Theorem 11.3 does not hold for rank one groups. We refer to [FG23, Example 6.8] for an example of a discrete SRS in $\text{SL}_{2}({\mathbb{R}})$ which is not an IRS, and note that such examples can be constructed in any rank one simple group.
@@
Let $\Gamma$ be a finitely generated group. Consider the deformation space $\text{Hom}(\Gamma,G)$ of homeomorphisms from $\Gamma$ to $G$ with the point-wise convergence topology. A map $\psi\in\text{Hom}(\Gamma,G)$ is called locally rigid if there is a neighborhood of $\psi$ which consists of conjugates. That is if $\phi\in\text{Hom}(\Gamma,G)$ is sufficiently close to $\psi$ then there is $g\in G$ such that $\phi=\psi^{g}$. A subgroup $\Gamma$ of $G$ is called locally rigid if the inclusion $i:\Gamma\hookrightarrow G$ is locally rigid. The classical local rigidity theorem due to Selbeg, Weil and Margulis, says the following:
@@
Let $G$ be a non-compact simple Lie group not locally isomorphic to $\text{SL}_{2}({\mathbb{R}})$ or $\text{SL}_{2}({\mathbb{C}})$. The every lattice in $G$ is locally rigid.
@@
It follows that all the relations $w\in R$ are satisfied on the set $\{\sigma_{n}:\sigma\in\Sigma\}$ and hence the assignment $\sigma\mapsto\sigma_{n}$ an homomorphism of $\Gamma$ which is close to the inclusion. By local rigidity we deduce that $\Lambda^{g_{n}}$ contains a conjugate of $\Gamma$. It follows that $\Lambda$ is a lattice, in contrast to the assumption. This completes the proof of Theorem 11.2. ∎
Summery of the proof:
@@
Let $G=G_{1}\times\cdots\times G_{n}$ be a connected centre-free semisimple Lie group without compact factors and with simple factors $G_{i},\leavevmode\nobreak\ i=1,\ldots,n$. Let $\nu$ be an ergodic discrete invariant random subgroup in $G$. Then $G$ decomposes to a product of semisimple factors $G=H_{1}\times\ldots\times H_{k}$ with $1\leq k\leq n$, such that almost surely the projection of a random subgroup to each $H_{i}$ is discrete while the projection of the intersection with $H_{i}$ to each proper factor of $H_{i}$ is dense.
@@
Let $G=G_{1}\times\cdots\times G_{n}$ be a connected centre-free semisimple Lie group such that each $G_{i}$ is simple of rank at least $2$. Let $\Lambda$ be a discrete subgroup. Then $\Lambda$ is confined if and only if there is a nontrivial semisimple factor $H\lhd G$ such that $\Lambda\cap H$ is a lattice in $H$. If $\Lambda$ is not confined then $\frac{1}{n}\sum_{i=0}^{n-1}\mu_{G}^{(i)}*\delta_{\Lambda}\to\delta_{\{1\}}$.
@@ 12 Spectral gap for products and a strong version of the NST
The impact of this theorem on the field is enormous. The proof is quite remarkable, both in the general idea and in the details. Margulis deduce’s the finiteness of $\Gamma/N$ from the tension between two analytic properties that cannot get along with each other for infinite groups, namely amenability and property $(T)$.
Recall that a group $H$ is amenable if the regular representation on $L_{2}(H)$ admits asymptotically invariant vectors, while property $(T)$ is a spectral gap with respect to all unitary representations and in particular the regular representation. Therefore if $H$ is both amenable and has $(T)$ then $L_{2}(H)$ must admit the constant non-zero functions which implies that $H$ is compact, or finite if it is also discrete.
@@
There are several works that generalizes and improves the normal subgroup theorem. For instance Thereom 5.1 can be regarded as a measure theoretic counterpart of the NST, while Theorem 11.1 can be regarded as a geometric variant of the NST and both are much stronger than the NST when they hold. However these works, as well as all other published generalizations of the NST for classical semisimple groups relied on Kazhdan’s property (T) and do not cover the general higher rank case.
@@
I will now review a joint work with Arie Levit and Uri Bader [BGL24] that will be posted on arXiv in a couple of weeks. We have managed to prove the following:
@@
In view of the ergodic decomposition of probability measure preserving actions, the lemma clearly reduces to the ergodic case, so let us assume that $\nu$ is $G_{1}$-ergodic. Using the large stabilizers of the action and Kakutani’s ergodic theorem for random walks we will prove that such measure must be also $G_{2}$-invariant.
@@
In a sense, Theorem 12.4 is established by proving a carful suitable approximated version of Lemma 12.5. That is, we establish that certain measures which are sufficiently $G_{1}$ invariant must also be $G_{2}$ almost invariant. However, the precise formulation and the proof are considerably more complicated and we refer the reader to [BGL24]. Below we shall outline how data this type implies the desired statements.
Since simple groups with rank at least $2$ have property $(T)$, Theorem 12.4 (as well as Theorem 12.2) follows from [FG23] in that case. Therefore we may suppose that $G$ has more than one factors. Let $\Lambda\leq G$ be an irreducibly confined subgroup. Throughout most of the proof we suppose that $\Lambda$ is also co-amenable. That is, there is a an asymptotically invariant sequence of unit vectors $f_{n}\in L_{2}(G/\Lambda)$. This means that if $\Omega$ is a compact set in $G$ then
@@
The above sketch is an over-simplification of the actual proof avoiding various nuances. For instance, it is crucial for us the the sequence $h_{n}$ is uniform in the sense that
@@
The presentation of the previous two subsection may make the impression that the ‘irreducibly confined’ assumption is only required in the reduction to the co-amenable case, and that for co-amenable groups a weaker irreducibility assumption might be sufficient. However, this is not the case. Even in the co-amenable case the ‘irreducibly confined’ assumption is crucial in our argument. The reason is that we apply a process of changing the probability measures and we have limited control on the way this is done. Therefore we need to assume some irreducibility also for Chabauty limits of stabalizers. The spectral gap result that we obtain in this case is quite general however:

### v2

@@ 10 Stationary random subgroups
The notion of invariant random subgroups has proven very effective to the study of certain types of problems. However, the restriction to ‘invariant’ measures limits the scope of problems that one can investigate, in particular as the groups we are studying are non-amenable. For example, the rigidity theorem of Stuck and Zimmer says that for higher rank semisimple Lie groups (with property $(T)$) all IRS comes from lattices. In particular one cannot hope to study thin subgroups or general discrete subgroups which are non-lattices using the theory of IRS. Also for simple groups of rank one, it is known that there are discrete subgroups that cannot be in the support of an IRS. On the other hand, every discrete subgroup is associated with some stationary random subgroups (hereafter SRS) as we shall explain below.
@@
Let $G$ be a connected semisimple Lie group. Let $\mu$ be a Borel regular probability measure on $G$ whose support generates a Zariski dense subgroup. Let $X$ be a compact $G$-space. A probability measure $\nu\in\text{Prob}(X)$ is called $\mu$-stationary if $\mu*\nu=\nu$ where
@@
We let $(B,\nu)$ denote the Poisson boundary of $G$ associated to $\mu$. Then $\nu$ is $\mu$-stationary and has the following universal property. Recall that on a compact convex space every probability measure has a unique barycenter. In particular, this applies to probability measures on the compact convex space $\text{Prob}(X)$.
@@
Let $X$ be a $G$-space. A probability measure $\nu_{0}$ on $X$ is $\mu$-stationary iff it is the barycenter of the measure $f_{*}\nu$ for some $G$-equivariant measurable map $f:B\to\text{Prob}(X)$. Moreover, the map $f$ is unique.
@@
Let $K$ be a fixed maximal compact subgroup of $G$. Below we will only consider measures $\mu$ which are bi-$K$-invariant. For such measures the Poisson boundary is $(G/P,\nu)$ where $P$ is a minimal parabolic subgroup and $\nu$ is the unique $K$-invariant measure on $G/P$ (the uniqueness is due to the fact that $K$ acts transitively on $G/P$ which is a consequence of the Iwasawa decomposition) . By the observation above we may vary the measure $\mu$ within this class (according to our needs) without affecting the property of being stationary.
@@
Motivated by [Ge18b] and the original paper of Kazhdan and Margulis [KM68] we have established in [GLM22] the following result which is a strong effective version of the Kazhdan–Margulis theorem. It is obviously stronger than the result from [Ge18b] as it is effective and applies to all SRS rather than IRS. Moreover, since it applies to all SRS it implies the Kazhdan–Margulis theorem for all discrete subgroups.
@@
$\nu(\{\Lambda\>:\>\Lambda\cap\mathrm{B}(r)\neq\{1\})\leq\beta r^{\delta}\quad\forall r<\rho.$ (4)
As a straightforward consequence, we obtain:
@@ 11 Confined subgroups have finite co-volume
In this section, we will explain the proof of the following result from [FG23] which confirmed a conjecture of Margulis.
@@
Geometrically, the theorem says that every random walk on a locally $X$-manifold of infinite volume spends most of the time in the $r$-thick part for every $r>0$.
@@
The aim is to prove that $\nu_{n}\to\delta_{\langle 1\rangle}$. Since $\text{Prob}(\text{Sub}(G))$ (with the weak-* topology) is compact, it is enough to prove that $\delta_{\langle 1\rangle}$ is the only limit of $\nu_{n}$. Thus let $\nu_{\infty}$ be a week-* limit of $\nu_{n}$, and we shall argue to show that $\nu_{\infty}=\delta_{\langle 1\rangle}$. In view of Theorem 10.5 $\nu_{\infty}$ is a discrete SRS, i.e. a $\nu_{\infty}$-random subgroup is almost surely discrete.
@@
Fix a minimal parabolic subgroup $P\leq G$ and recall that the Poisson boundary of $(G,\mu)$ is $(G/P,\nu_{P})$ where $\nu_{P}$ is the unique $K$-invariant probability measure on $G/P$. Let $\nu_{0}$ be an SRS. In view of Furstenberg’s theorem (see 10.1) $\nu_{0}$ is the barycentre of a $G$-equivariant measurable map $\phi:G/P\to\text{Prob}(\text{Sub}(G))$, that is:
@@
Let $G$ be a higher rank simple Lie group. Let $\mu$ be a smooth probability measure on $G$ and let $(X,\nu)$ be a probability $\mu$-stationary action of $G$. Then either $\nu$ is $G$-invariant, or there exists a proper parabolic subgroup $Q\subset G$ and a measure preserving $G$-equivariant map999 By measure preserving we only mean that $\nu(\pi^{-1}(A))=\mu_{Q}(A)$ for every Borel set $A\subset G/Q$. In particular, we do not require that it is a relative measure preserving extension. $\pi\colon(X,\nu)\to(G/Q,\mu_{Q}),$ where $\mu_{Q}$ is the unique $\mu$-stationary measure on $G/Q$.
@@
Arguing by way of contradiction, suppose that $\nu_{0}$ is a non-invariant discrete SRS. It follows from Theorem 11.4 that there is a parabolic subgroup $Q$ containing $P$ such that for $\nu_{P}$ almost all $gP\in G/P$ the measure $\phi(gP)$ is a $gPg^{-1}$-invariant discrete random subgroup of $gQg^{-1}$. The contradiction, thus follows from the following:
@@
Let $P\leq Q$ be a parabolic subgroup of $G$. The only $P$-invariant discrete random subgroup of $Q$ is $\delta_{\langle 1\rangle}$.
@@
Let $Q$ be a parabolic subgroup of $G$. Let $Q=LN$ be a Levi decomposition of $Q$ and let $A_{L}$ be the center of $L$. Then, any discrete $A_{L}$-invariant random subgroup of $Q$ is contained in $L$ almost surely.
@@
Let $\lambda$ be a discrete $A_{L}$-invariant random subgroup of $Q$. Suppose by way of contradiction that $\lambda(\{\Gamma\subset L\})<1$. Then there is a bounded identity neighborhood $V_{2}$ in $Q$ such that $\lambda(\Omega)>0$ where $\Omega:=\{\Gamma:\Gamma\cap V_{2}\setminus L\neq\emptyset\}$. Furthermore, we may take $V_{2}$ to be of the form $V_{2}=W\cdot U_{2}$ where $W\subset L$ and $U_{2}\subset N$ and suppose that they are preserved by $a^{-1}$ (e.g. we can suppose that $\log U_{2}$ is a norm ball in the Lie algebra of $N$). Since $\lambda$ is discrete we can chose a small identity neighborhood $U_{1}\subset N$ so that setting $V_{1}=W\cdot U_{1}$ we have
@@
We remark that the analog of stiffness Theorem 11.3 does not hold for rank one groups. We refer to [FG23, Example 6.8] for an example of a discrete SRS in $\text{SL}_{2}({\mathbb{R}})$ which is not an IRS. Relying on [O24] one can construct examples of that nature in any Gromov hyperbolic group and hence in any rank one simple Lie group.
@@
Let $\Gamma$ be a finitely generated group. Consider the deformation space $\text{Hom}(\Gamma,G)$ of homeomorphisms from $\Gamma$ to $G$ with the point-wise convergence topology. A map $\psi\in\text{Hom}(\Gamma,G)$ is called locally rigid if there is a neighborhood of $\psi$ which consists of conjugates. That is if $\phi\in\text{Hom}(\Gamma,G)$ is sufficiently close to $\psi$ then there is $g\in G$ such that $\phi=\psi^{g}$. A subgroup $\Gamma$ of $G$ is called locally rigid if the inclusion $i:\Gamma\hookrightarrow G$ is locally rigid. The classical local rigidity theorem due to Selbeg, Weil, and Margulis, says the following:
@@
Let $G$ be a non-compact simple Lie group not locally isomorphic to $\text{SL}_{2}({\mathbb{R}})$ or $\text{SL}_{2}({\mathbb{C}})$. Then every lattice in $G$ is locally rigid.
@@
It follows that all the relations $w\in R$ are satisfied on the set $\{\sigma_{n}:\sigma\in\Sigma\}$ and hence the assignment $\sigma\mapsto\sigma_{n}$ a homomorphism of $\Gamma$ which is close to the inclusion. By local rigidity, we deduce that $\Lambda^{g_{n}}$ contains a conjugate of $\Gamma$. It follows that $\Lambda$ is a lattice, in contrast to the assumption. This completes the proof of Theorem 11.2. ∎
Summary of the proof:
@@
Let $G=G_{1}\times\cdots\times G_{n}$ be a connected center-free semisimple Lie group without compact factors and with simple factors $G_{i},\leavevmode\nobreak\ i=1,\ldots,n$. Let $\nu$ be an ergodic discrete invariant random subgroup in $G$. Then $G$ decomposes to a product of semisimple factors $G=H_{1}\times\ldots\times H_{k}$ with $1\leq k\leq n$, such that almost surely the projection of a random subgroup to each $H_{i}$ is discrete while the projection of the intersection with $H_{i}$ to each proper factor of $H_{i}$ is dense.
@@
Let $G=G_{1}\times\cdots\times G_{n}$ be a connected center-free semisimple Lie group such that each $G_{i}$ is simple of rank at least $2$. Let $\Lambda$ be a discrete subgroup. Then $\Lambda$ is confined if and only if there is a nontrivial semisimple factor $H\lhd G$ such that $\Lambda\cap H$ is a lattice in $H$. If $\Lambda$ is not confined then $\frac{1}{n}\sum_{i=0}^{n-1}\mu_{G}^{(i)}*\delta_{\Lambda}\to\delta_{\{1\}}$.
@@ 12 Spectral gap for products and a strong version of the NST
The impact of this theorem on the field is enormous. The proof is quite remarkable, both in the general idea and in the details. Margulis deduces the finiteness of $\Gamma/N$ from the tension between two analytic properties that cannot get along with each other for infinite groups, namely amenability and property $(T)$.
Recall that a group $H$ is amenable if the regular representation on $L^{2}(H)$ admits asymptotically invariant vectors, while property $(T)$ is a spectral gap with respect to all unitary representations and in particular the regular representation. Therefore if $H$ is both amenable and has $(T)$ then $L^{2}(H)$ must admit the constant non-zero functions which implies that $H$ is compact, or finite if it is also discrete.
@@
There are several works that generalize and improve the normal subgroup theorem. For instance Thereom 5.1 can be regarded as a measure-theoretic counterpart of the NST, while Theorem 11.1 can be regarded as a geometric variant of the NST and both are much stronger than the NST when they hold. However these works, as well as all other published generalizations of the NST for classical semisimple groups relied on Kazhdan’s property (T) and do not cover the general higher rank case.
@@
I will now review a very recent joint work with Arie Levit and Uri Bader [BGL24]. We have managed to prove the following:
@@
In view of the ergodic decomposition of probability measure preserving actions, the lemma clearly reduces to the ergodic case, so let us assume that $\nu$ is $G_{1}$-ergodic. Using the large stabilizers of the action and Kakutani’s ergodic theorem for random walks we will prove that such a measure must be also $G_{2}$-invariant.
@@
In a sense, Theorem 12.4 is established by proving a careful suitable approximated version of Lemma 12.5. That is, we establish that certain measures which are sufficiently $G_{1}$ invariant must also be $G_{2}$ almost invariant. However, the precise formulation and the proof are considerably more complicated and we refer the reader to [BGL24]. Below we shall outline how data of this type implies the desired statements.
Since simple groups with rank at least $2$ have property $(T)$, Theorem 12.4 (as well as Theorem 12.2) follows from [FG23] in that case. Therefore we may suppose that $G$ has more than one factor. Let $\Lambda\leq G$ be an irreducibly confined subgroup. Throughout most of the proof, we suppose that $\Lambda$ is also co-amenable. That is, there is an asymptotically invariant sequence of unit vectors $f_{n}\in L_{2}(G/\Lambda)$. This means that if $\Omega$ is a compact set in $G$ then
@@
The above sketch is an oversimplification of the actual proof avoiding various nuances. For instance, it is crucial for us the the sequence $h_{n}$ is uniform in the sense that
@@
The presentation of the previous two subsections may make the impression that the ‘irreducibly confined’ assumption is only required in the reduction to the co-amenable case and that for co-amenable groups a weaker irreducibility assumption might be sufficient. However, this is not the case. Even in the co-amenable case, the ‘irreducibly confined’ assumption is crucial in our argument. The reason is that we apply a process of changing the probability measures and we have limited control on the way this is done. Therefore we need to assume some irreducibility also for Chabauty limits of stabilizers. The spectral gap result that we obtain in this case is quite general, however: